Hi, I have to dertermine the speed distribution of the fluid (vz(r)) in the same configuration of the exercice 4.88 of the book Fluid Mechanics of Frank M White. But instead of a fixed outside cylinder, the outside cylinder is moving to its left with a constant speed V. For what ratio value V/U, the shear force on the two walls of the cylinder will be equal?
Answer #1


flow is steady b 12 2 Tag 2 flow is fully axial ie Vecco e Vo=0 Vz Pure OL, SA ie Vz= f(01) only using スー momentum egy +t? 12 Nz = (U+V) enx + U - (U+v) en a V+ (U+V) in en 19 in un (9) r=b b Shear stress Trab) = u dvz u (U+V). I en 19) de luzb b Shea